# Quartic soliton solutions of a normal dispersion based mode-locked laser

### v1

@@ III.4 Energy and Width
One point that has been raised about quartic solitons in mode-locked lasers is their ability to reach very high energy and very short temporal widths since their conservative versions have energy that scales inversely with the width cubed. In a previous work [7], we have shown that this scaling was only valid for this model for some of the solitons that exist for negative 4OD. To understand how energy and width depend on the parameters $(g_{0},T_{2})$, we used all the stable MAS and HAS that were obtained to produce Figs. 11 and 11 showing energy ($E$) versus width ($\tau$). Each type of solution presents a different behavior. For HAS, the logarithms of energy and width fall in a single line giving the relation $E=14.5\,\tau^{2.00}$ (for energies in pJ and widths in ps) as presented in Fig. 11. This energy-width relation is not in agreement with the linear energy-width expression obeyed by the approximated profile given by Eq. (10). The reason for this disagreement should be related to the definition of the half maximum. The approximate profile does not predict the thin top spike of the HAS, such that, the half maximum of the approximate profile can be significantly lower than the actual one and then its full width can be significantly higher than the actual width. Regarding the energy-width relation for MAS, there are different curves for each $g_{0}$ and each curve does not follow a power law. Moreover, the energy may decrease or increase with the width depending on the region of the parameter space, such that, for lower $T_{2}$ the energy tends do decrease with width but, for higher $T_{2}$, the energy tends to increase with width. These results for MAS do not follow the $E\propto\tau^{-3}$ that was observed for the CGLE solutions in [16].
Figure 11: Energy versus temporal width for 11 MAS and several $g_{0}$ (m-1) values as indicated in the figure and for all the obtained HAS solutions 11 in base 10 logarithms. The line in 11 is a linear fit to the points whose equation is also on the graph.
@@ IV Modulation instability
Figure 13: Evolution of a perturbed homogeneous solution for 13 $g_{0}=1.46$ and $T_{2}=80$, 13 for $g_{0}=1.46$ and $T_{2}=120$.The final line in each graph corresponds to the final $|q|$ output, being at the same scale as the 3D plot in the case of 13 but multiplied by 10 in the case of 13 in order to become visible.
@@ V Effect of second and third order dispersion
Considering $T_{2}$ and $g_{0}$ inside the region of existence and stability of HAS, namely $T_{2}=60$ fs and $g_{0}=1.4$ m-1, we observed the evolution of the corresponding HAS in case $\beta_{2}$ is nonzero. The evolution tends to new stationary profiles that are shown in Fig. 14 for $\beta_{2}=\pm 0.1$ ps2m-1. These results show that positive $\beta_{2}$ increases the amplitude and energy of the pulse, eliminating the thin spike of the HAS. On the other hand, negative $\beta_{2}$ decreases the amplitude and energy and increases the structure of the top spike. A similar approach was used to understand the effect of $\beta_{3}$. The final stationary soliton profiles are shown in Fig. 14 side by side with the pure quartic HAS for the same parameters as in the above paragraph but with $\beta_{2}=0$ and $\beta_{3}=\pm 0.05$ ps3m-1, which is close to the maximum allowed $\beta_{3}$ that preserves stationary evolution of the pulse. The profiles are asymmetrical and have slightly higher amplitude than the pure quartic HAS. In the figure, the peak position of the nonzero $\beta_{3}$ profiles are to the left and right of the HAS peak, depending on the sign of $\beta_{3}$, which arose from the displacement from the original location of the pulse due to the nonzero velocity that $\beta_{3}$ imposes on the pulses. Note that the asymmetry and velocity are identical but have opposite directions depending on the sign of $\beta_{3}$.
Figure 14: Power profiles of the stationary solitons in the region of stable HAS ($T_{2}=60$ fs and $g_{0}=1.4$ m-1) for 14 $\beta_{2}=\pm 0.1$  ps2m-1 and 14 $\beta_{3}=\pm 0.05$  ps3m-1 as indicated in the legends. The profile for $\beta_{2}=0$ is added for comparison. A zoom of the peaks is shown, in 14 for both cases of nonzero $\beta_{2}$.
@@
Figure 16: Power profiles of the stationary solitons in the region of stable MAS ($T_{2}=150$ fs and $g_{0}=1.48$ m-1) for 16 $\beta_{2}=\pm 0.1$  ps2m-1 and 16 for $\beta_{3}=\pm 0.1$  ps3m-1 as indicated in the legends. The profile for $\beta_{2}=0$ is added for comparison.
@@ VI Conclusions
An in-depth study about soliton solutions of a distributed model for a mode-locked laser, dominated by positive fourth order dispersion, was carried out. Some of the types of solutions have no practical importance since they are always unstable. Two types of solutions exist and are stable for a relatively large regions of linear gain and spectral filtering parameters. They even coexist and are stable in a particular region of parameters, thus, the model presents effects of bistability, which could have application in optical logic gates. These two kinds of solutions differ in their energy-width relation. The energy of the high amplitude solution scales with the square of the FWHM measured from the thin spike on the top of the pulse that is characteristic of this type of solution. On the other hand, the energy of the medium amplitude solution has no fixed relation with the temporal width, it may increase or decrease with the temporal width. For lower spectral filtering and higher linear gain parameters, this type of solution has the advantage of being energetic and short. Moreover, this same type of solution, i.e, the MAS may be self excited through modulational instability which corresponds to an advantage for the development of self-starting pulsed lasers. Both MAS and HAS do survive in the presence of second and third order dispersion contributions. The preliminary results also reveal that these lower order dispersion contributions together with 4OD may give rise to new types of solutions that will be worth studying in future works. All these results should be interesting to both theoretical and experimental communities studying and implementing mode-locked lasers. They will be useful for finding the best parameters of the laser that will produce pulses with the desired characteristics.
@@ Acknowledgements.
This work has received funding from the Fundação para a Ciência e a Tecnologia (FCT) within the Projects LA/P/0037/2020, UIDB/50014/2020, UIDB/50025/2020 and UIDP/50025/2020. D. Malheiro acknowledges a PhD fellowship from FCT with code 2024.01557.BD.

### v2

@@ III.4 Energy and Width
One point that has been raised about quartic solitons in mode-locked lasers is their ability to reach very high energy and very short temporal widths since their conservative versions have energy that scales inversely with the width cubed. In a previous work [7], we have shown that this scaling was only valid for this model for some of the solitons that exist for negative 4OD. To understand how energy and width depend on the parameters $(g_{0},T_{2})$, we used all the stable MAS and HAS that were obtained to produce Figs. 12 and 12 showing energy ($E$) versus width ($\tau$). Each type of solution presents a different behavior. For HAS, the logarithms of energy and width fall in a single line giving the relation $E=16\,\tau^{2.0}$ (for energies in pJ and widths in ps) as presented in Fig. 12. This energy-width relation is in agreement with the one estimated using the approximated profile given by Eq. (10). Regarding the energy-width relation for MAS, there are different curves for each $g_{0}$ and each curve does not follow a power law. Moreover, the energy may decrease or increase with the width depending on the region of the parameter space, such that, for lower $T_{2}$ the energy tends do decrease with width but, for higher $T_{2}$, the energy tends to increase with width. These results for MAS do not follow the $E\propto\tau^{-3}$ that was observed for the CGLE solutions in [16].
Figure 12: Energy versus temporal width for 12 MAS and several $g_{0}$ (m-1) values as indicated in the figure and for all the obtained HAS solutions 12 in base 10 logarithms. The line in 12 is a linear fit to the points whose equation is also on the graph.
@@ IV Modulation instability
Figure 14: Evolution of a perturbed homogeneous solution for 14 $g_{0}=1.46$ m-1 and $T_{2}=80$ fs, 14 for $g_{0}=1.46$ m-1 and $T_{2}=120$ fs.The final line in each graph corresponds to the final $|q|$ output, being at the same scale as the 3D plot in the case of 14 but multiplied by 10 in the case of 14 in order to become visible.
@@ V Effect of second and third order dispersion
Considering $T_{2}$ and $g_{0}$ inside the region of existence and stability of HAS, namely $T_{2}=60$ fs and $g_{0}=1.4$ m-1, we observed the evolution of the corresponding HAS in case $\beta_{2}$ is nonzero. The evolution tends to new stationary profiles that are shown in Fig. 15 for $\beta_{2}=\pm 0.1$ ps2m-1. These results show that positive $\beta_{2}$ increases the amplitude and energy of the pulse, eliminating the thin spike of the HAS. On the other hand, negative $\beta_{2}$ decreases the amplitude and energy and increases the structure of the top spike. A similar approach was used to understand the effect of $\beta_{3}$. The final stationary soliton profiles are shown in Fig. 15 side by side with the pure quartic HAS for the same parameters as in the above paragraph but with $\beta_{2}=0$ and $\beta_{3}=\pm 0.05$ ps3m-1, which is close to the maximum allowed $\beta_{3}$ that preserves stationary evolution of the pulse. The profiles are asymmetrical and have slightly higher amplitude than the pure quartic HAS. In the figure, the peak position of the nonzero $\beta_{3}$ profiles are to the left and right of the HAS peak, depending on the sign of $\beta_{3}$. In the presence of $\beta_{3}$, the pulse deforms asymmetrically and the effective group velocity differs from the assumed group velocity calculated for the central wavelength. Thus, in the time referential used in equation (1) that travels with the latter velocity, the pulses are displaced from the initial temporal position by $t(z)=t(0)+vz$, with $v$ being the shift in velocity introduced by $\beta_{3}$. Note that the asymmetry and velocity are identical but have opposite directions depending on the sign of $\beta_{3}$.
Figure 15: Power profiles of the stationary solitons in the region of stable HAS ($T_{2}=60$ fs and $g_{0}=1.4$ m-1) for 15 $\beta_{2}=\pm 0.1$  ps2m-1 and 15 $\beta_{3}=\pm 0.05$  ps3m-1 as indicated in the legends. The profile for $\beta_{2}=0$ is added for comparison. A zoom of the peaks is shown, in 15 for both cases of nonzero $\beta_{2}$.
@@
Figure 17: Power profiles of the stationary solitons in the region of stable MAS ($T_{2}=150$ fs and $g_{0}=1.48$ m-1) for 17 $\beta_{2}=\pm 0.1$  ps2m-1 and 17 for $\beta_{3}=\pm 0.1$  ps3m-1 as indicated in the legends. The profile for $\beta_{2}=0$ is added for comparison.
@@ VI Conclusions
An in-depth study about soliton solutions of a distributed model for a mode-locked laser, dominated by positive fourth order dispersion, was carried out. Some of the types of solutions have no practical importance since they are always unstable. Two types of solutions exist and are stable for a relatively large regions of linear gain and spectral filtering parameters. They even coexist and are stable in a particular region of parameters, thus, the model presents effects of bistability, which could have application in optical logic gates. These two kinds of solutions differ in their energy-width relation. The energy of the high amplitude solution scales with the square of the FWHM measured from the thin spike on the top of the pulse that is characteristic of this type of solution. On the other hand, the energy of the medium amplitude solution has no fixed relation with the temporal width, it may increase or decrease with the temporal width. For lower spectral filtering and higher linear gain parameters, this type of solution has the advantage of being energetic and short. Moreover, this same type of solution, i.e, the MAS may be self excited through modulational instability which corresponds to an advantage for the development of self-starting pulsed lasers. Both MAS and HAS do survive in the presence of second and third order dispersion contributions. The preliminary results also reveal that these lower order dispersion contributions together with 4OD may give rise to new types of solutions that will be worth studying in future works. All these results should be interesting to both theoretical and experimental communities studying and implementing mode-locked lasers. In particular, they will be useful for identifying laser implementations that produce pulses with the desired characteristics—either by selecting setups with similar parameters to those used here, or by using the adimensionalized variables and parameters, also introduced here, to predict pulse behavior for other achievable configurations.
@@ Acknowledgements.
This work has received funding from the Fundação para a Ciência e a Tecnologia (FCT) within the Projects LA/P/0037/2020, UIDB/50014/2020, UIDB/50025/2020, UIDP/50025/2020, and PTDC/FIS-OUT/3882/2020. D. Malheiro acknowledges a PhD fellowship from FCT with code 2024.01557.BD.